I have an unknown organic compound containing carbon, hydrogen and oxygen. Combustion analysis of 4.05 g sample of the unknown produced 8.07 g CO2 and 3.76 g H2O. The molar mass of the unknown was determined to be 180 +/- 4 g/mol. What is the molecular formula of the unknown?
Answer #1

From the given information first we have to find number of moles of C, H, o # weight of co- 8.079. moles of con- 8.018 = 0.18# Mass of oxygen = (mass of sample)-(Mass of C+H) = 4.058 - (2-1969 + 0.4189) = 4.059-2.614 g = 1.436 g Moles of o= 1.436 g =